# Representation Ensembling for Synergistic Lifelong Learning with Quasilinear Complexity

### v18

@@ E.1 EWC
$=\tilde{\mathcal{O}}(nT).$ (10)
E.2 Total Replay
@@
$=\tilde{\mathcal{O}}(nT)$ (11)
E.3 ProgNN
@@
$=\tilde{\mathcal{O}}(nT)$ (12)
Appendix F Simulated Results
In each simulation, we constructed an environment with two tasks. For each, we sample 750 times from the first task, followed by 750 times from the second task. These 1,500 samples comprise the training data. We sample another 1,000 hold out samples to evaluate the algorithms. We fit a random forest (RF) (technically, an uncertainty forest which is an honest forest with a finite-sample correction [79]) and a SynF. For SynN, we have used a deep network (DN) architecture with two hidden layers each having $10$ nodes. Similarly, for SynN experiments we did $100$ repetitions and reported the results after smoothing it using moving average with a window size of $5$. For the SynF experiments we used $1000$ repetitions and reported the mean of these repetitions. We repeat this process 30 times to obtain errorbars. Error bars in all cases were negligible.
F.1 Gaussian XOR
@@
Figure 2: Top: 750 samples from 3 spirals (left) and 5 spirals (right). Bottom left: SynF outperforms RF on 3 spirals when 5 spirals data is available, demonstrating backward transfer in SynF. Bottom center: SynF outperforms RF on 5 spirals when 3 spirals data is available, demonstrating forward transfer in SynF. Bottom right: Transfer Efficiency of SynF. The forward (solid) and backward (dashed) curves are the ratio of the generalization error of SynF to RF in their respective figures. SynF demonstrates decreasing forward transfer and increasing backward transfer in this environment.
F.2 Spirals
A description of the distributions for the two tasks is as follows: let $K$ be the number of classes and $S\sim$ multinomial$(\frac{1}{K}\vec{1}_{K},n)$. Conditioned on $S$, each feature vector is parameterized by two variables, the radius $r$ and an angle $\theta$. For each sample, $r$ is sampled uniformly in $[0,1]$. Conditioned on a particular class, the angles are evenly spaced between $\frac{4\pi(k-1)t_{K}}{K}$ and $\frac{4\pi(k)t_{K}}{K}$ where $t_{K}$ controls the number of turns in the spiral. To inject noise along the spiral, we add Gaussian noise to the evenly spaced angles $\theta^{\prime}:\theta=\theta^{\prime}+\mathcal{N}(0,\sigma_{K}^{2})$. The observed feature vector is then $(r\;\cos(\theta),r\;\sin(\theta))$. In Figure 2 we set $t_{3}=2.5$, $t_{5}=3.5$, $\sigma_{3}^{2}=3$ and $\sigma_{5}^{2}=1.876$.
Consider an environment with a three spiral and five spiral task (Figure 2). In this environment, axis-aligned splits are inefficient, because the optimal partitions are better approximated by irregular polytopes than by the orthotopes provided by axis-aligned splits. The three spiral data helps the five spiral performance because the optimal partitioning for these two tasks is relatively similar to one another, as indicated by positive forward transfer. This is despite the fact that the five spiral task requires more fine partitioning than the three spiral task. Because SynF grows relatively deep trees, it over-partitions space, thereby rendering tasks with more coarse optimal decision boundaries useful for tasks with more fine optimal decision boundaries. The five spiral data also improves the three spiral performance.
Appendix G Real Data Extended Results
@@
Table 5: Performance metric: average $\mathsf{Transfer}$ after $10$ tasks calculated for different algorithms on CIFAR 10x10 ($500$ samples per task).
Algorithms $\mathsf{Transfer}(\pm\text{std})$
SynN $\mathbf{~{}~{}0.19}(\pm 0.04)$
SynF $\mathbf{~{}~{}0.13}(\pm 0.02)$
Model Zoo ${}~{}0.09(\pm 0.04)$
ProgNN ${}~{}0.11(\pm 0.09)$
LMC $-0.05(\pm 0.04)$
CoSCL $-0.06(\pm 0.07)$
DF-CNN $-0.11(\pm 0.08)$
EWC $-0.15(\pm 0.04)$
Total Replay $-0.15(\pm 0.03)$
Partial Replay $-0.13(\pm 0.03)$
SynF(resource constrained) $\mathbf{~{}~{}0.05}(\pm 0.03)$
LwF $-0.05(\pm 0.03)$
O-EWC $-0.14(\pm 0.04)$
SI $-0.16(\pm 0.03)$
ER $-0.13(\pm 0.12)$
A-GEM $-0.19(\pm 0.14)$
TAG $-0.32(\pm 0.04)$
None $-0.24(\pm 0.10)$
Table 6: Performance metric: average $\mathsf{Transfer}$ after $10$ tasks calculated for different algorithms on 5-dataset.
Algorithms $\mathsf{Transfer}(\pm\text{std})$
SynN $\mathbf{-0.27}(\pm 0.22)$
SynF $\mathbf{-0.05}(\pm 0.11)$
Model Zoo ${}~{}0.24(\pm 0.12)$
EWC $-1.06(\pm 0.60)$
Total Replay $-0.18(\pm 0.22)$
Partial Replay $-0.27(\pm 0.26)$
LwF $-0.39(\pm 0.48)$
O-EWC $-1.07(\pm 0.60)$
SI $-1.15(\pm 0.69)$
ER $-0.78(\pm 1.03)$
A-GEM $-0.55(\pm 0.90)$
TAG $-0.56(\pm 0.58)$
None $-1.15(\pm 1.30)$
Table 7: Performance metric: average $\mathsf{Transfer}$ after $10$ tasks calculated for different algorithms on Split Mini-Imagenet.
Algorithms $\mathsf{Transfer}(\pm\text{std})$
SynN $\mathbf{~{}~{}0.02}(\pm 0.10)$
SynF $\mathbf{~{}~{}0.10}(\pm 0.04)$
Model Zoo ${}~{}0.23(\pm 0.10)$
EWC $-0.29(\pm 0.12)$
Total Replay ${}~{}0.06(\pm 0.13)$
Partial Replay ${}~{}0.00(\pm 0.10)$
LwF ${}~{}0.02(\pm 0.08)$
O-EWC $-0.21(\pm 0.10)$
SI $-0.14(\pm 0.12)$
ER $-0.02(\pm 0.27)$
A-GEM ${}~{}0.06(\pm 0.26)$
TAG $-0.05(\pm 0.15)$
None $-0.22(\pm 0.23)$
Table 8: Performance metric: average $\mathsf{Transfer}$ after $10$ tasks calculated for different algorithms on FOOD1k 50X20.
Algorithms $\mathsf{Transfer}(\pm\text{std})$
SynN $\mathbf{~{}~{}0.31}(\pm 0.06)$
SynF $\mathbf{~{}~{}0.14}(\pm 0.04)$
Model Zoo ${}~{}0.13(\pm 0.08)$
LwF $-0.06(\pm 0.13)$
Table 9: Hyperparameters for SynF in spoken digit experiment.
@@
G.1 CIFAR 10x10
Table 10: Task splits for CIFAR 10x10.
Task # Image Classes
1 apple, aquarium fish, baby, bear, beaver, bed, bee, beetle, bicycle, bottle
2 bowl, boy, bridge, bus, butterfly, camel, can, castle, caterpillar
3 chair, chimpanzee, clock, cloud, cockroach, couch, crab, crocodile, cup, dinosaur
4 dolphin, elephant, flatfish, forest, fox, girl, hamster, house, kangaroo, keyboard
5 lamp, lawn mower, leopard, lion, lizard, lobster, man, maple tree, motor cycle, mountain
6 mouse, mushroom, oak tree, orange, orchid, otter, palm tree, pear, pickup truck, pine tree
7 plain, plate, poppy, porcupine, possum, rabbit, raccoon, ray, road, rocket
8 rose, sea, seal, shark, shrew, skunk, skyscraper, snail, snke, spider
9 squirrel, streetcar, sunflower, sweet pepper, table, tank, telephone, television, tiger, tractor
10 train, trout, tulip, turtle, wardrobe, whale, willow tree, wolf, woman, worm
G.2 CIFAR 10x10 Repeated Classes
We also considered the setting where each task is defined by a random sampling of 10 out of 100 classes with replacement. This environment is designed to demonstrate the effect of tasks with shared subtasks, which is a common property of real world lifelong learning tasks. Supplementary Figure 3 shows transfer of SynF and SynN on Task 1.
Figure 3: SynF and SynN transfer knowledge effectively when tasks share common classes. Each task is a random selection of 10 out of the 100 CIFAR-100 classes. Both SynF and SynN demonstrate monotonically increasing transfer efficiency for up to 20 tasks.
@@ G.2 CIFAR 10x10 Repeated Classes
Supplementary Table 10 shows the image classes associated with each task number.
G.3 Spoken Digit experiment

### v19

@@ A.3 Training Time Complexity Analysis
(13) $=\tilde{\mathcal{O}}(nT).$
Total Replay
@@
(14) $=\tilde{\mathcal{O}}(nT)$
ProgNN
@@
(15) $=\tilde{\mathcal{O}}(nT)$
A.4 Simulated Results
In each simulation, we constructed an environment with two tasks. For each, we sample 750 times from the first task, followed by 750 times from the second task. These 1,500 samples comprise the training data. We sample another 1,000 hold out samples to evaluate the algorithms. For SiLLy-N, we have used a deep network (DN) architecture with two hidden layers each having $10$ nodes. Similarly, for SiLLy-N experiments we did $100$ repetitions and reported the results after smoothing it using moving average with a window size of $5$.
Gaussian XOR
@@ A.4 Simulated Results
A.5 Real Data Extended Results
@@
Table 4: Hyperparameters for SiLLy-F in spoken digit experiment.
@@
CIFAR 10x10 Repeated Classes
We also considered the setting where each task is defined by a random sampling of 10 out of 100 classes with replacement. This environment is designed to demonstrate the effect of tasks with shared subtasks, which is a common property of real world lifelong learning tasks. Supplementary Figure 2 shows transfer of SiLLy-F and SiLLy-N on Task 1.
Figure 2: SiLLy-N and SiLLy-F transfer knowledge effectively when tasks share common classes. Each task is a random selection of 10 out of the 100 CIFAR-100 classes. Both SiLLy-F and SiLLy-N demonstrate monotonically increasing transfer efficiency for up to 20 tasks.
@@
Spoken Digit experiment